# Learning to Optimize for Mixed-Integer Non-linear Programming

### v1

@@ 1 Introduction
  • We propose two novel differentiable correction layers that effectively handle the non-differentiability of integer outputs in deep learning models. Combined with a penalty method for soft constraint satisfaction, we are able to learn a neural network mapping from instance parameters to solutions through gradient-based learning. Not only does this yield an extremely fast heuristic at test time, but the method is also self-supervised and thus efficiently trainable.
@@ 5 Learning to Optimize MINLPs with Correction Layers
5.1 Rounding Classification
@@
This approach can be viewed as an end-to-end learnable version of the Relaxation Enforced Neighborhood Search (RENS) algorithm (Berthold, 2014). Instead of explicitly searching the neighborhood of the relaxed solution, the neural network implicitly learns the corrections required to achieve a feasible integer solution by exploring the solution space near the integer variables while updating the continuous variables.
@@
5.2 Learnable Threshold
Here, the slope is set to 10 to sharpen the Sigmoid function.
@@ Methods.
The experiments evaluate our learning-based methods, Rounding Classification (RC) and Learnable Threshold (LT), against traditional exact optimization (EX), which can compute optimal solutions but is often computationally expensive, and heuristic-based approaches such as Rounding after Relaxation (RR) and root node solutions (N1), which offer faster results without quality guarantees. Rounding after Learning (RL) is also used as a baseline, applying rounding to the relaxed solution generated by learning-based methods rather than learning-to-round as our correction layers do. Note that baselines EX and N1 include a wide range of heuristics that are embedded in the MINLP solver of choice (Gurobi or SCIP) and that are executed in conjunction with the tree search procedure; we are also implicitly comparing to these heuristics, not just to the exact search. As such, the competing methods cover a broad spectrum of optimization strategies, from exact solvers to fast heuristics, allowing for a comprehensive evaluation of solution quality and computational efficiency. Table 1 provides an overview of the methods. A 60-second time limit is enforced for all methods and problems.
@@
Rounding after Learning* RL* Learning-based method that learns from the relaxed problem, then rounds the continuous solution to the nearest integer.
@@
Root Node Solution N1 A feasible solution obtained at the root node using heuristics after continuous relaxation with cutting planes.
@@
We tested the methods on a variety of optimization problems, including integer convex quadratic problems, simple integer non-convex problems, and high-dimensional mixed-integer Rosenbrock problems. These problem classes were selected to cover both convex and non-convex scenarios and to evaluate the scalability of the methods in higher-dimensional settings. Each method was assessed in terms of objective value, constraint violation, and solving time, providing a comprehensive view of their performance across different types of problems. In addition, we performed experiments on integer linear programs in Appendix D; these experiments simply serve to show that our methods could be applied to the integer linear case, though the use of MILP solvers may be preferable.
@@ 6.1 Convex Quadratic Problem
where the coefficients $\bm{Q}\in\mathbb{R}^{n\times n}$, $\bm{p}\in\mathbb{R}^{n}$, and $\bm{A}\in\mathbb{R}^{m\times n}$ were fixed and generated using random values, while $\bm{b}\in\mathbb{R}^{m}$ were treated as parametric coefficients (input features), varying across instances. Further details on the generation of $\bm{Q},\bm{p},\bm{A},$ and $\bm{b}$ can be found in Section C.1.
@@
We experimented with quadratic problems of different sizes, from $5$ decision variables and $5$ constraints ($5\times 5$) up to $500\times 500$. The results in Table 2 summarize the performance of all methods across different problem sizes. The RC and LT methods exhibit robust performance across the board, achieving objective values second only to EX while consistently maintaining low percentages of infeasible solutions and fast solution times across all problem sizes. These methods achieve several orders of magnitude speed-ups, scaling effectively even for large instances up to $500\times 500$. The exact solver EX, while performing well on smaller problem sizes, fails to produce any solutions for instances of size $100\times 100$ and larger within the 60-second time limit, highlighting its limitations when handling more complex problems. N1, on the other hand, can find feasible solutions within a short time frame for smaller cases but suffers from severe numerical instability as the problem size increases. When scaled to $200\times 200$, N1 also fails to produce a solution. Both the RR and RL methods, which rely on rounding relaxations of either the solver or the prediction model, encounter significant feasibility challenges. Overall, this analysis underscores that learning-based methods like RC and LT offer considerable advantages in both solution quality and computational speed, especially for large-scale problems, compared to exact solvers or other heuristics.
@@
In addition to evaluating solution quality, feasibility, and solving/inference times, we also measured the offline training times for our two approaches on different problem sizes. These results, along with training times for other problem types, are presented in Appendix E, where it is evident that the training times for the learning-based methods scale well with problem size.
@@ 6.2 Simple Non-convex Problem
where the sine function is applied element-wise to the decision variables $\bm{x}$. This introduces non-convexity into the problem, making it more challenging compared to the convex case. The coefficients $\bm{Q}$, $\bm{p}$, $\bm{A}$, and $\bm{b}$ are generated in the same way as in the quadratic formulation (Section C.1). In addition, the scales of the problem and the experiment setting are also identical to those of the quadratic problems.
@@ 6.3 Multi-Dimensional Mixed-Integer Rosenbrock Problem
where $\bm{x}\in\mathbb{R}^{n}$ are continuous decision variables and $\bm{y}\in\mathbb{Z}^{n}$ are integer decision variables. The vectors $\bm{p}\in\mathbb{R}^{n}$ and $\bm{q}\in\mathbb{R}^{n}$ are fixed for each instance, while the parameters $b$ and $\bm{a}$ vary. Details on how the parameters $\bm{p},\bm{q},\bm{a},$ and $b$ were generated can be found in Section C.2.
We conducted experiments on mixed-integer Rosenbrock problems with the number of decision variables ranging from 2 to 20,000; the number of constraints was fixed at 5. The results in Table 4 show that RC and LT exhibit strong performance, even outperforming EX in smaller cases. However, as the problem size increases to 10,000 variables, a noticeable decline in feasibility is observed for both RC and LT, while solver-based methods such as EX, N1, and RR fail to produce any solutions. As seen in previous experiments, RR and RL, which rely on rounding relaxations, continue to suffer from significant infeasibility issues.
@@ 7 Conclusion
We have introduced a new learning-based heuristic method for MINLP. Our approach includes two novel correction layers—rounding classification and learnable threshold—that enable neural networks to generate high-quality integer solutions while preserving gradient information for training through backpropagation. These layers allow us to tackle optimization tasks with discrete variables and non-linear constraints in a way that is scalable and computationally efficient.
@@ Appendix B Neural Network Structure and Hyperparameters
Appendix C Detailed Problem Setup and Parameter Sampling
C.1 Convex Quadratic and Simple Non-convex Problems
@@ C.2 Ronsenbrock Problems
Appendix D Experiments on Integer Linear Programs
@@
The ILP instances were generated with random coefficients for the objective function and constraint matrices in a manner consistent with the setup for the convex quadratic problems. Each instance featured varying right-hand side values, which were treated as input features. More details on the parameter generation process can be found in Appendix C.1.
@@
Appendix E Training Time Comparison
@@
Tables 6, 7, 8 and 9, summarize the training times (in seconds) required by each method for problems of different scales. These results offer a clear view of the computational demands as problem complexity grows.

### v2

@@ 1 Introduction
  • We propose two novel differentiable correction layers that effectively handle the non-differentiability of integer outputs in deep learning models. Combined with a penalty method for soft constraint satisfaction, we are able to learn a neural network mapping from instance parameters to solutions through gradient-based learning. Not only does this yield an extremely fast heuristic at test time, but the method is also self-supervised and thus efficiently trainable. Additionally, we introduce a strategy of progressively increasing penalties to further improve feasibility in the generated solutions.
@@ 5 Learning to Optimize MINLPs with Correction Layers
These approaches can be viewed as an end-to-end learnable version of the Relaxation Enforced Neighborhood Search (RENS) algorithm (Berthold, 2014). Instead of explicitly searching the neighborhood of the relaxed solution, the neural network implicitly learns the corrections required to achieve a feasible integer solution by exploring the solution space near the integer variables while updating the continuous variables.
Rounding Classification.
@@
Learnable Threshold.
Here, the slope is set to 10 to sharpen the Sigmoid function.
@@
Ablation Study.
The two baselines involve only some parts of our full pipeline and thus serve as ablations. Rounding after Learning (RL) is used as a baseline, applying rounding to the relaxed solution generated by learning-based methods rather than learning-to-round as our correction layers do. In Rounding with STE (RS), the continuous values predicted by the first neural network $\pi_{\Theta_{1}}$ are directly rounded without any learned correction mechanism, i.e., without using the correction network $\varphi_{\Theta_{2}}$; STE is used to backpropagate through the rounding operator. For further details, refer to Appendix C.
@@ Methods.
Table 1 provides an overview of all the methods used in the following experiments. A 60-second time limit is enforced for all methods and problems. The experiments evaluate our learning-based methods, Rounding Classification (RC) and Learnable Threshold (LT), against traditional exact optimization (EX), which can compute optimal solutions but is often computationally expensive, and heuristic-based approaches such as Rounding after Relaxation (RR) and root node solutions (N1), which offer faster results without quality guarantees. Note that baselines EX and N1 include a wide range of heuristics that are embedded in the MINLP solver of choice (Gurobi or SCIP) and that are executed in conjunction with the tree search procedure; we are also implicitly comparing to these heuristics, not just to the exact search. As such, the competing methods cover a broad spectrum of optimization strategies, from exact solvers to fast heuristics, allowing for a comprehensive evaluation of solution quality and computational efficiency. In addition, two ablation baselines are included: Rounding after Learning (RL), which applies direct rounding to the relaxed solution generated by the learning-based method without any corrective adjustments, and Rounding with STE (RS), which uses the Straight-Through Estimator to handle non-differentiable rounding but without a learning-based correction layer.
@@
Rounding after Learning* RL* Learning-based method that learns from the relaxed problem, then rounds the inferred solution to the nearest integer.
Rounding with STE* RS* Learning-based method where the rounding process is differentiable with STE but does not involve learning.
@@
Root Node Solution N1 A feasible solution obtained at the root node, which uses heuristics after continuous relaxation with cutting planes.
@@
We tested the methods on a variety of optimization problems, including integer convex quadratic problems, simple integer non-convex problems, and high-dimensional mixed-integer Rosenbrock problems. These problem classes were selected to cover both convex and non-convex scenarios and to evaluate the scalability of the methods in higher-dimensional settings. Each method was assessed in terms of objective value, constraint violation, and solving time, providing a comprehensive view of their performance across different types of problems. In addition, we performed experiments on integer linear programs in Appendix E; these experiments simply serve to show that our methods could be applied to the integer linear case, though the use of MILP solvers may be preferable.
@@ 6.1 Convex Quadratic Problem
RS Obj Mean $0.941$ $-1.097$ $-3.945$ $-11.982$ $-11.003$ $-24.598$ $-57.312$
Obj Median $0.777$ $-1.409$ $-3.956$ $-12.002$ $-11.205$ $-24.566$ $-57.312$
% Infeasible $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$
Time (Sec) $0.0014$ $0.0018$ $0.0020$ $0.0015$ $0.0020$ $0.0043$ $0.0038$
where the coefficients $\bm{Q}\in\mathbb{R}^{n\times n}$, $\bm{p}\in\mathbb{R}^{n}$, and $\bm{A}\in\mathbb{R}^{m\times n}$ were fixed and generated using random values, while $\bm{b}\in\mathbb{R}^{m}$ were treated as parametric coefficients (input features), varying across instances. Further details on the generation of $\bm{Q},\bm{p},\bm{A},$ and $\bm{b}$ can be found in Section D.1.
We experimented with quadratic problems of different sizes, from $5$ decision variables and $5$ constraints ($5\times 5$) up to $500\times 500$. The results in Table 2 summarize the performance of all methods across different problem sizes. The RC and LT methods exhibit robust performance across the board, achieving objective values second only to EX while consistently maintaining low percentages of infeasible solutions and fast solution times across all problem sizes. These methods achieve several orders of magnitude speed-ups, scaling effectively even for large instances up to $500\times 500$. Without learning-based correction, the objective values of RS are generally worse than those of RC and LT, but there are no infeasible solutions. The exact solver EX, while performing well on smaller problem sizes, fails to produce any solutions for instances of size $100\times 100$ and larger within the 60-second time limit, highlighting its limitations when handling more complex problems. N1, on the other hand, can find feasible solutions within a short time frame for smaller cases but suffers from severe numerical instability as the problem size increases. When scaled to $200\times 200$, N1 also fails to produce a solution. Both the RR and RL methods, which rely on rounding relaxations of either the solver or the prediction model, encounter significant feasibility challenges. Overall, this analysis underscores that learning-based methods like RC and LT offer considerable advantages in both solution quality and computational speed, especially for large-scale problems, compared to exact solvers or other heuristics.
@@
In addition to evaluating solution quality, feasibility, and solving/inference times, we also measured the offline training times for our two approaches on different problem sizes. These results, along with training times for other problem types, are presented in Appendix F, where it is evident that the training times for the learning-based methods scale well with problem size.
@@ 6.2 Simple Non-convex Problem
where the sine function is applied element-wise to the decision variables $\bm{x}$. This introduces non-convexity into the problem, making it more challenging compared to the convex case. The coefficients $\bm{Q}$, $\bm{p}$, $\bm{A}$, and $\bm{b}$ are generated in the same way as in the quadratic formulation (Section D.1). In addition, the scales of the problem and the experiment setting are also identical to those of the quadratic problems.
@@ 6.3 Multi-Dimensional Mixed-Integer Rosenbrock Problem
where $\bm{x}\in\mathbb{R}^{n}$ are continuous decision variables and $\bm{y}\in\mathbb{Z}^{n}$ are integer decision variables. The vectors $\bm{p}\in\mathbb{R}^{n}$ and $\bm{q}\in\mathbb{R}^{n}$ are fixed for each instance, while the parameters $b$ and $\bm{a}$ vary. Details on how the parameters $\bm{p},\bm{q},\bm{a},$ and $b$ were generated can be found in Section D.2.
We conducted experiments on mixed-integer Rosenbrock problems with the number of decision variables ranging from 2 to 20,000; the number of constraints was fixed at 5. The results in Table 4 show that RC and LT exhibit strong performance, even outperforming EX in smaller cases, while RS performs slightly worse than RC and LT in terms of objective value. However, as the problem size increases to 10,000 variables, a noticeable decline in feasibility is observed for both RC and LT, while solver-based methods such as EX, N1, and RR fail to produce any solutions. As seen in previous experiments, RR and RL, which rely on rounding relaxations, continue to suffer from significant infeasibility issues.
@@
RS Obj Mean $25.07$ $65.38$ $634.95$ $6566.85$ $72904.02$
Obj Median $25.28$ $63.86$ $619.46$ $6441.98$ $72533.57$
% Infeasible $0\%$ $2\%$ $0\%$ $0\%$ $35\%$
Time (Sec) $0.0018$ $0.0016$ $0.0017$ $0.0027$ $0.0066$
6.4 Effect of Increasing Penalty Parameters
In this section, we evaluate the effect of progressively increasing penalty parameters on the above benchmark problems. This experiment is inspired by penalty methods commonly used in optimization, such as the Augmented Lagrangian Method (Hestenes, 1969). The weight of the penalty term is typically increased throughout the iterations. This gradual increase in penalty helps avoid numerical difficulties caused by overly strict constraints at the early stages while enhancing constraint satisfaction as the solution converges.
Table 5: Results for Increasing Penalty Parameters on Benchmark Problems. See the caption of Table 2 for details.
Problem Size Method Obj Mean Obj Median % Infeasible Time (Sec)
Convex Quadratic 100×100 RC $-10.660$ $-10.697$ $1\%$ $0.0036$
LT $-11.548$ $-11.617$ $0\%$ $0.0036$
200×200 RC $-25.350$ $-25.412$ $0\%$ $0.0033$
LT $-25.110$ $-25.297$ $0\%$ $0.0033$
500×500 RC $-60.040$ $-60.405$ $4\%$ $0.0062$
LT $-57.507$ $-57.590$ $1\%$ $0.0063$
Simple Non-Convex 100×100 RC $1.773$ $1.718$ $0\%$ $0.0038$
LT $1.932$ $1.917$ $0\%$ $0.0035$
200×200 RC $1.662$ $1.584$ $0\%$ $0.0037$
LT $2.163$ $2.060$ $0\%$ $0.0034$
[…]
Rosenbrock 200×5 RC $519.85$ $467.59$ $0\%$ $0.0034$
LT $618.51$ $624.10$ $0\%$ $0.0031$
2000×5 RC $7276.96$ $7460.26$ $0\%$ $0.0042$
LT $5780.85$ $5590.29$ $0\%$ $0.0044$
20000×5 RC $63626.35$ $58977.61$ $29\%$ $0.0044$
LT $63734.97$ $64431.54$ $39\%$ $0.0044$
In this experiment, we initialized the penalty weight at $100$, with a growth rate of $1.05$. Table 5 presents the results under this approach. In general, we observe that the method successfully improves the feasibility at the cost of objective value. It reflects the trade-off between enhancing stricter constraint satisfaction and achieving better solutions. However, for the 20000×5 Rosenbrock problem, even with progressively increasing penalties, the predictor still yields many infeasible solutions.
@@ 7 Conclusion
We have introduced a new learning-based heuristic method for MINLP. Our approach includes two novel correction layers—rounding classification and learnable threshold—that enable neural networks to generate high-quality integer solutions while preserving gradient information for training through backpropagation. These layers allow us to tackle optimization tasks with discrete variables and non-linear constraints in a way that is scalable and computationally efficient. As a self-supervised approach, our method does not require collecting optimal solutions as labels, significantly reducing the time and effort typically needed for data collection. Additionally, we proposed a strategy of progressively increasing penalties to further enhance the feasibility of the predicted solutions.
@@ Appendix B Neural Network Structure and Hyperparameters
Appendix C Details for Ablation Study
C.1 Rounding after Learning
Rounding after Learning serves as an ablation study, where direct rounding is applied to the relaxed solution generated by the learning-based model $\pi_{\Theta_{1}}(\cdot)$. In this case, the loss function minimizes the objective based on the continuous relaxed solution $\bar{\bf x}\in\mathbb{R}^{n_{r}+n_{z}}$. However, when direct rounding is applied, the resulting integer solution can cause significant deviations in the objective value and may fail to satisfy feasibility constraints. This highlights the importance of end-to-end learning, where the model receives updates directly from the ultimate loss function.
C.2 Rounding with Straight-through Estimator
As another part of our ablation study, we then implement the STE to handle the non-differentiable rounding operations: 1) In this approach, the relaxed solution $\bar{\bf x}^{i}_{z}$ for integer variables is rounded down, and the STE technique enables effective gradients to pass through the rounding step during backpropagation. 2) To determine the rounding direction ${\bf b}$, we employ the Gumbel-Sigmoid function to provide approximation gradients for backward passes and output a binarized result during the forward passes.
Algorithm 2 Rounding with STE for Learning-to-optimize MINLPs: Forward Pass.
1:Training instance ${{\bm{\xi}}^{i}}$ and neural networks $\pi_{\Theta_{1}}(\cdot)$
3:Round integer variables down: $\hat{{\bf x}}^{i}_{z}\leftarrow\lfloor\bar{{\bf x}}^{i}_{z}\rfloor$
[…]
5:Update integer variables: $\hat{{\bf x}}^{i}_{z}\leftarrow\hat{{\bf x}}^{i}_{z}+{\bf b}^{i}$
6:return $\hat{\bf x}^{i}$
While this method does not involve any learning mechanism for rounding correction $\varphi$, the model in this ablation setup directly rounds to the nearest integer without leveraging information from the whole relaxed solution and problem coefficients. This helps to measure how learning-based rounding can refine the rounding process and enhance overall performance.
Appendix D Detailed Problem Setup and Parameter Sampling
D.1 Convex Quadratic and Simple Non-convex Problems
@@ D.2 Ronsenbrock Problems
Appendix E Experiments on Integer Linear Programs
@@
The ILP instances were generated with random coefficients for the objective function and constraint matrices in a manner consistent with the setup for the convex quadratic problems. Each instance featured varying right-hand side values, which were treated as input features. More details on the parameter generation process can be found in Appendix D.1.
@@
Appendix F Training Time Comparison
@@
Tables 7, 8, 9 and 10, summarize the training times (in seconds) required by each method for problems of different scales. These results offer a clear view of the computational demands as problem complexity grows.